The proportion of the maximum allowed time to complete a project is modeled as a beta...

  1. The proportion of the maximum allowed time to complete a project is modeled as a beta distribution with α=2 and β=3 . What is the probability that the proportion of the maximum time exceeds 0.6 ?
Answer #1
f(x)=(1/B(α,β))*xα-1*(1-x)β-1 = 12*x*(1-x)2   for 0<X<1

P(X>0.6) = f(x) dx = 12(x+x3-2x2) dx =12*(x2/2+x4/4-2x3/3)|10.6 =1-0.8208 =0.1792